OpenStudy (anonymous):

Solve 5x2 = -30x - 65.

OpenStudy (anonymous):

\[5x^2 = -30x - 65 \] rewrite as \[5x^2+30x+65=0\] divide by 5 \[x^2+6x+13=0\] and do from there

OpenStudy (anonymous):

x = -3 ± 2i x = -3 ± 4i x = -6 ± 2i x = -6 ± 4i

OpenStudy (anonymous):

guess it it one of those

OpenStudy (anonymous):

still confused @satellite73

OpenStudy (anonymous):

\[x^2+6x+13=0\] is what you need to solve do you know how to do it"

OpenStudy (kva1992):

cant you just use the snowflake method?

OpenStudy (anonymous):

lol if i had ever heard of snowflake what is that, walt disney math?

OpenStudy (kva1992):

haha well i call it the diamond method but i recently heard it referred to as the snowflake method

OpenStudy (kva1992):

actually you have to do the quadratic nvm lol

OpenStudy (anonymous):

not heard of either must be new math in any case lets solve it \[x^2+6x+13=0\\ x^2+6x=-13\\ (x+3)^2=-13+9=-4\\ x+3=\sqrt{-4}\\x+3=\pm2i\\ x=-3\pm2i\]

OpenStudy (kva1992):



OpenStudy (kva1992):

thats the diamond i was taught as diamond im sure its referred to as something different

OpenStudy (anonymous):

i have absolutely no clue as to what you are doing

OpenStudy (kva1992):

i was just showing you what the diamond is lol usually used for factoring polynomials instead of the guessing method

OpenStudy (anonymous):

thank you @satellite73

OpenStudy (anonymous):

yw